OpenStudy (anonymous):

Determine if (1/3,-5) is a solution to 3x-10y=51. Please show all of your work.

OpenStudy (bahrom7893):

plug in 1/3 instead of x

OpenStudy (bahrom7893):

and -5 instead of y and see if u get 51

OpenStudy (bahrom7893):

3*(1/3)-10*(-5)=1+50=51, so yes it is a solution